OpenStudy (trisha001):

Will fan and medal for help. Thanks. The dimension of a rectangle can be given by x+7 and x+2. If the area of the rectangle is 66 square inches, what are the dimensions of the rectangle?

OpenStudy (isaidavila):

A=bh A=(x+7)(x+2) 66=(x+7)(x+2) Solve for that

OpenStudy (trisha001):

Wait could you help me out? Am I supposed to solve both of those or? I am new at this and I'm really confused. I have no idea what I am doing I'm sorry

OpenStudy (isaidavila):

Just find values for x 66=6*11 so X+7=11 X=4 Got it?

OpenStudy (trisha001):

Yeah I think I understand now thank you.

OpenStudy (isaidavila):

You can check the answer replacing 4 for x in the equation above

OpenStudy (trisha001):

wait is it a=x2+9x+14?

OpenStudy (trisha001):

@isaidavila

OpenStudy (isaidavila):

Yeah